### v2

# GS-ROR: 3D Gaussian Splatting for Reflective Object Relighting via SDF Priors

Abstract.
3D Gaussian Splatting (3DGS) has shown a powerful capability for novel view synthesis due to its detailed expressive ability and highly efficient rendering speed. Unfortunately, creating relightable 3D assets with 3DGS is still problematic, particularly for reflective objects, as its discontinuous representation raises difficulties in constraining geometries. Inspired by previous works, the signed distance field (SDF) can serve as an effective way for geometry regularization. However, a direct incorporation between Gaussians and SDF significantly slows training. To this end, we propose GS-ROR for reflective objects relighting with 3DGS aided by SDF priors. At the core of our method is the mutual supervision of the depth and normal between deferred Gaussians and SDF, which avoids the expensive volume rendering of SDF. Thanks to this mutual supervision, the learned deferred Gaussians are well-constrained with a minimal time cost. As the Gaussians are rendered in a deferred shading mode, while the alpha-blended Gaussians are smooth, individual Gaussians may still be outliers, yielding floater artifacts. Therefore, we further introduce an SDF-aware pruning strategy to remove Gaussian outliers, which are located distant from the surface defined by SDF, avoiding the floater issue. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods in terms of relighting quality. Our method also exhibits competitive relighting quality compared to NeRF-based methods with at most 25% of training time and allows rendering at 200+ frames per second on an RTX4090.
neural rendering, Gaussian splatting, relighting
Corresponding author.
submissionid: 360journal: TOGjournal: TOGccs: Computing methodologies Rendering
Figure 1. We present an SDF-aided Gaussian Splatting framework for Reflective Object Relighting (GS-ROR) from multi-view images. We show five relighting results with reflective highlights (left) and their normal estimation (right), including Potion and Tbell from NeRO (Liu et al., 2023), Helmet and Toaster from Ref-NeRF (Verbin et al., 2022), and Qilin from NeILF++ (Zhang et al., 2023), where Qilin is from the real scene. Our method demonstrates a robust geometry reconstruction for reflective surfaces and faithful material decomposition, leading to photo-realistic and real-time reflective object relighting.
1. Introduction
Creating relightable 3D assets from multi-view images has been a long-standing and challenging task in computer graphics and vision, as the decomposition of lighting, materials, and geometries is highly ill-posed. Particularly, the decomposition becomes more difficult for reflective objects, as their appearances are highly view-dependent, and a minor error on the surface leads to a significant difference. Existing approaches (Liu et al., 2023; Li et al., 2024) have shown impressive relighting quality for reflective objects by leveraging the neural radiance field (NeRF) and the signed distance field (SDF). Unfortunately, these methods require a long training and rendering time. In this paper, we aim at reflective objects relighting given multi-view images, simultaneously achieving high-quality relighting and short training/rendering time costs.
Most recently, Kerbl et al. (2023) proposed 3D Gaussian Splatting (3DGS), boosting the rendering speed significantly and achieving more detailed appearance modeling. A concurrent work (Ye et al., 2024) has introduced deferred Gaussian splatting for reflective objects, significantly improving the novel view synthesis (NVS) quality. Despite the impressive NVS quality, the relighting with 3DGS becomes problematic due to the Gaussians’ discontinuity and high flexibility, especially for reflective objects. Extensive efforts have been made to improve the relighting quality by regularizing the geometry in terms of the normal and depth (Gao et al., 2024; Jiang et al., 2024; Liang et al., 2024). However, they still suffer from erroneous surfaces and floater artifacts for reflective objects.
In this paper, we propose a novel framework GS-ROR for reflective objects relighting by leveraging SDF priors in the deferred Gaussian splatting. A direct incorporation between Gaussians and SDF significantly slows training, due to the expensive computational cost of SDF rendering. For this, we propose to supervise the deferred Gaussians and SDF mutually in our framework, so-called mutual-supervision, which allows constraining deferred Gaussians with SDFs while avoiding SDF rendering. Specifically, the depth/normal of deferred Gaussians is supervised by SDF for regularization, while the SDF is supervised by Gaussians via the depth/normal rather than SDF rendering. This way, the learned deferred Gaussians are better constrained with a minimal computational cost. However, we still notice that Gaussians have floater artifacts occasionally, as the deferred Gaussians (with alpha blending) rather than individual Gaussians are constrained by SDF. Therefore, we introduce an SDF-aware pruning strategy to remove Gaussian outliers to avoid this floater issue. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods in terms of relighting quality. Meanwhile, it exhibits competitive relighting quality compared to NeRF-based methods with at most 25% of training time. It also allows real-time rendering of reflective objects with 200+ frames per second on an RTX4090. To summarize, our main contributions include
  • a GS-ROR framework for reflective object relighting, achieving high-quality relighting while maintaining short training time and allowing real-time rendering,
  • a mutual supervision approach between deferred Gaussians and the SDF, producing robust geometry and avoiding local minima with minimal computational increase, and
  • an SDF-aware Gaussian pruning strategy, which prunes outlier Gaussian, avoiding the float artifacts.
2. Related Work
2.1. Neural representations for multi-view stereo
Since the presence of NeRF (Mildenhall et al., 2020), numerous neural implicit representations (Fridovich-Keil et al., 2022; Barron et al., 2022; Müller et al., 2022; Chen et al., 2022) have been proposed and gained remarkable progress in the field of multi-view stereo. These NeRF-based models adopt Multi-Layer Perceptrons (MLPs) (Barron et al., 2021, 2022; Verbin et al., 2022) or grid-like representations (Müller et al., 2022; Chen et al., 2022; Fridovich-Keil et al., 2022) to represent geometry and view-dependent appearance, optimized via volume rendering. Subsequently, NeuS (Wang et al., 2021) links the SDF with the density in NeRF and enables the optimization of surfaces with volume rendering. With extra regularization, these SDF-based models (Li et al., 2023; Rosu and Behnke, 2023; Wang et al., 2023) achieve detailed surface reconstruction from RGB images. Owing to carefully designed architectures and specific optimization, some methods (Müller et al., 2022; Fridovich-Keil et al., 2022) enable fast training and real-time radiance field rendering.
3DGS (2023) sheds light on the multi-view stereo in the rasterization framework, providing impressive real-time NVS results. 3DGS utilizes the discrete explicit 3D Gaussian primitive and projects these primitives onto the image plane to obtain the pixel color via alpha blending, which avoids the consuming sampling operation. However, 3DGS imposes no geometry constraint while optimizing, leading to noisy depth and normal, which hinders the usage of 3DGS in many downstream tasks. To resolve this issue, existing methods construct different constraints on Gaussians by introducing 2D Gaussians (Huang et al., 2024) or Gaussian Surfel (Dai et al., 2024). SuGaR (Guédon and Lepetit, 2024) extracts a mesh from the Gaussians and binds the flattened Gaussians to the surface to optimize further. These methods improve the geometry quality significantly, which is orthogonal to our relighting framework and can further improve our relighting quality. We leave it for future work.
Several previous works introduce SDF into 3DGS. NeuSG (Chen et al., 2023) introduces SDF into 3DGS to improve the details of surface reconstruction by optimizing NeuS and 3DGS jointly. Despite the high quality of the reconstructed surface, the training time is tens of times longer than 3DGS. GSDF (Yu et al., 2024) utilizes the depth from Gaussian to guide sampling for SDF, but they still need to render color in the SDF branch, leading to higher efficiency. While our method also incorporates the SDF and Gaussians, the key difference is that it drops SDF rendering, improving the geometry quality with a low time cost.
2.2. Inverse rendering
Inverse rendering aims to decompose geometry, material, and light from multi-view RGB images. Most existing methods adopt the geometry formulation from NeRF in either density or SDF manner, namely NeRF-based method. In terms of material modeling, most methods utilize the Disney Principled BRDF model (McAuley et al., 2013), while NeRFactor (Zhang et al., 2021b) pretrains a BRDF MLP with a measured BRDF dataset. Considering the light modeling, early works (Zhang et al., 2021a; Boss et al., 2021b; Bi et al., 2020; Boss et al., 2021a; Zhang et al., 2022a) consider direct light only, and more recent works (Srinivasan et al., 2021; Zhang et al., 2022b; Jin et al., 2023; Zhang et al., 2023; Yang et al., 2023) model indirect lighting and visibility, leading to a higher-quality decomposition. NeRF Emitter (Ling et al., 2024) incorporates NeRF as a non-distant environment emitter and uses importance sampling of NeRF to reduce rendering variance. However, these NeRF-based methods suffer from slow rendering speed and high memory usage in training.
More recently, Gaussian-based inverse rendering methods have begun to emerge. These methods regularize Gaussians to obtain better geometry in various ways. Gaussian Shader (Jiang et al., 2024) and GIR (Shi et al., 2023) link the shortest axis of Gaussian with the normal. The Gaussian shader, Relightable Gaussian (Gao et al., 2024), and GS-IR (Liang et al., 2024) supervise the Gaussian normal with the depth normal. To model the indirect illumination, Relightable Gaussian, GIR, and GS-IR compute the visibility via one-step ray tracing and represent the indirect illumination with SH coefficients. All these Gaussian-based relighting methods cannot relight reflective objects, while our method enables such scenarios.
2.3. Reflective objects NVS and relighting
As a special and challenging case, NVS and relighting of reflective objects have drawn much attention. Ref-NeRF (Verbin et al., 2022) replaces the outgoing radiance with the incoming radiance and models the shading explicitly with an integrated direction encoding (IDE) to improve the quality of reflective materials. MS-NeRF (Yin et al., 2023) designs a multi-space neural representation to model the reflection of mirror-like planes. Spec-NeRF (Ma et al., 2024) introduces a learnable Gaussian directional encoding to better model view-dependent effects. Spec-Gaussian (Yang et al., 2024) uses anisotropic spherical Gaussian to model view-dependent highlights. 3DGS-DR (Ye et al., 2024), as a concurrent work, introduces deferred shading into 3DGS. We also utilize a deferred framework, but for relighting, and together with other key components.
Another group of methods enables relighting for reflective materials. NeRO (Liu et al., 2023) proposes a novel light representation consisting of two MLPs for direct and indirect light, respectively. With explicit incorporation of the rendering equation, NeRO reconstructs reflective objects of high quality. Wang et al. (2024) represent the indirect light with a 5-dimensional function, so-called the 5D neural plenoptic function. Thanks to material-aware integral positional encoding, their method reduces the light sample per pixel during training and provides more accurate material decomposition. TensoSDF (Li et al., 2024) designs a roughness-aware incorporation between the radiance and reflectance field for robust geometry reconstruction and proposes a novel tensorial representation with SDF, enabling fast SDF query. Although their representation is faster than the original SDF, it still needs hours of training time and cannot achieve real-time rendering. DeferredGS (Wu et al., 2024) utilizes the idea of deferred shading to improve the quality of the reflective object relighting in the Gaussian framework. However, their method needs to distill a reasonable geometry from a pretrained SDF, taking more than 3 hours and thus losing the efficiency in training. Our method takes a further step in the reflective object relighting with 3D Gaussians and provides high-quality relighting results while keeping highly efficient training and rendering.
Figure 2. The geometry from Gaussian is under-constrained and thus erroneous, while it is much better after utilizing the priors from the SDF.
Figure 3. Overview of our framework. The architecture of our proposed method consists of two geometry representations (i.e., Gaussian primitive and TensoSDF). In the deferred Gaussian pipeline, the shading parameters (i.e., albedo a, roughness r and metallicity m), normal and depth are projected to the image plane and alpha blended. The pixel color $C_{\rm gs}$ is calculated using the split-sum approximation and supervised by ground truth color $C_{\rm gt}$. In the TensoSDF, we sample rays originated from camera center o and view direction v and query the SDF value and gradient for each point $p$ along the ray $\textbf{o}+t\textbf{v}$. The normal $\textbf{n}_{\rm sdf}$ and depth $D_{\rm sdf}$ are obtained via volume rendering, which is supervised mutually with the normal $\textbf{n}_{\rm gs}$ the depth $D_{\rm gs}$ from Gaussians. Note no color network is used in the SDF part, and only the geometry attributes are volume rendered.
3. Preliminaries and motivation
In this section, we briefly review two ways for 3D scene representations and then discuss our motivation.
3D Gaussian Splatting
3DGS represents a scene with a set of 3D Gaussians whose distribution is defined as
(1) $G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)},$
where $x$ is a position in the scene, $\mu$ is the mean of the Gaussian, and $\Sigma$ denotes the covariance matrix of the 3D Gaussian, which is factorized into a scaling matrix $S$ and a rotation matrix $R$ as $\Sigma=RSS^{T}R^{T}$. To render an image, 3DGS projects the 3D Gaussians onto the image plane and employs alpha blending on the sorted 2D Gaussians as
(2) $C=\sum^{n}_{i=0}{c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})},$
where $c_{i}$ is the color of each Gaussian, and $\alpha_{i}$ is given by evaluating a projected 2D Gaussian with covariance $\Sigma^{\prime}$ multiplied with a learned per-point opacity.
TensoSDF
An SDF expresses a scene by encoding the distance from the surface of an object, which can be expressed implicitly by an MLP (Wang et al., 2021) or in a hybrid way by combining explicit grids with a tiny MLP (Li et al., 2024), which allows more efficient training and rendering. Depth and normal are two significant attributes of geometry. In SDF, The point-wise depth $D_{i}$ is defined as the distance from the camera to the sample, and the point-wise normal $\textbf{n}_{i}$ is the gradient of SDF $\nabla_{x}$ w.r.t. position $x$ (Zhang et al., 2021a). Given $n$ points along the ray $p_{i}=\textbf{o}+t\textbf{v}$ where o is the camera origin and v is the view direction, we can aggregate these point-wise attributes to obtain the attributes (i.e., depth, normal) of the hit point of the surface as:
(3) $D_{\rm sdf}=\sum^{n}_{i=0}w_{i}D_{i},\textbf{n}_{\rm sdf}=\sum^{n}_{i=0}w_{i}\textbf{n}_{i}$
where $w_{i}$ is the weight of the $i$-th point derived from the SDF value.
SDF and NeRF can be unified by NeuS and VolSDF (Wang et al., 2021; Yariv et al., 2021), which maps an SDF value to a density distribution. The distribution in NeuS is defined as:
(4) $\phi_{s}(s)=\gamma e^{-\gamma s}/(1+e^{-\gamma s})^{2},$
where $s$ is the SDF value and $\gamma$ is the inverse of standard deviation. Note that $\gamma$ is a trainable parameter, and $1/\gamma$ approaches zero as the network training converges.
Among the SDF-based methods, TensoSDF utilizes a tensorial representation for SDF, consisting of a tensorial encoder and MLP decoder, which is formulated as follows:
(5) $V_{p}=v^{X}_{k}\circ M^{YZ}_{k}\oplus v^{Y}_{k}\circ M^{XZ}_{k}\oplus v^{Z}_{k}\circ M^{XY}_{k},s=\Theta(V_{p},p),$
where $v^{m}_{k}$ and $M^{\widetilde{m}}_{k}$ represent the $k$-th vector and matrix factors of their corresponding spatial axes $m$, and $\widetilde{m}$ denotes the two axes orthogonal to $m$ e.g., $\widetilde{X}=YZ$. $\circ$ and $\oplus$ represent the element-wise multiplication and concatenation operations. $V_{p}$ is the latent vector from the tensorial encoder and then is decoded with the position $p$ by a tiny MLP $\Theta$ to get the SDF value $s$. TensoSDF backbone enables a faster convergence speed and low-cost SDF query.
Motivation
The above two representations can both express a 3D scene, where the Gaussian representation is explicit and discontinuous, and the SDF is implicit (or hybrid) and continuous. These two representations lead to different rendering methods: splatting or rasterization for Gaussians and sphere tracing for SDF. While the Gaussian representation significantly boosts the rendering speed to a real-time level, it has difficulties constraining geometries, leading to inferior geometries, as shown in Fig. 2. A straightforward idea is to incorporate these two representations by leveraging the SDF as a prior. However, rendering SDF is time-consuming, which hinders the benefits of 3DGS. Therefore, the essential problem is an elegant incorporation between the Gaussians and the SDF.
4. Method
In this section, we present our framework GS-ROR for reflective object relighting. First, we present a deferred Gaussian splatting for relighting (Sec. 4.1). Then, we introduce the SDF priors into the deferred Gaussian splatting (Sec. 4.2) for geometry regularization. Finally, we propose an SDF-aware pruning strategy to refine the Gaussian distributions (Sec. 4.3).
4.1. Deferred Gaussian splatting for relighting
Figure 4. Two Gaussians with a minor normal difference are overlapped to model an opaque surface. In the forward rendering, the BRDF values are computed w.r.t. to their own normal and are then alpha blended to form the final rendering, which is equivalent to a broader BRDF lobe, leading to a blurry rendering, eventually. In contrast, in the deferred shading, the BRDF is computed w.r.t. the deferred normal, maintaining the sharpness of reflective objects.
Starting from GS-IR (Liang et al., 2024), we first construct a forward Gaussian splatting baseline for relighting, using the Disney Principled BDRF model (McAuley et al., 2013) and environment map to model the material and light. We drop the occlusion for efficiency and modify the normal formulation following Gaussian Shader (Jiang et al., 2024) as the combination of the shortest axis direction $u$ and an offset $n_{\delta}$ as $\textbf{n}=u+n_{\delta}$. The radiance from view v for each Gaussian is computed by the split-sum approximation, following previous work (Jiang et al., 2024; Liang et al., 2024), and then aggregated by alpha blending. With multi-view images as inputs, the reflectance parameters at each Gaussian can be learned, besides the opacity and the environment map, enabling relighting under a novel environment map.
However, this simple Gaussian-based relighting baseline produces blurry relighting results for the reflective surface. The main reason behind this phenomenon is the alpha blending of Gaussians. As shown in Fig. 4, the Gaussians are overlapped to model an opaque surface with slightly different normals. While this difference among Gaussians causes a negligible error for rough objects, it amplifies the BRDF lobe for specular objects, leading to a blurry rendering. The key to addressing this issue is to blend the normals of Gaussians first and then perform shading, i.e., deferred shading.
Specifically, we perform alpha blending on all the attributes (e.g., normal, albedo, roughness) onto the image plane:
(6) $F=\sum^{n}_{i=0}{f_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})},$
where $\alpha_{i},\alpha_{j}$ share the same definition as in Eqn. (2), and $f_{i}$ is a attribute of the $i$-th Gaussian. Then, we use split-sum approximation to compute the pixel-wise color.
The Gaussians are learned with the following loss function:
(7) $\mathcal{L}_{\rm gs}=\mathcal{L}_{\rm c}+\lambda_{\rm nd}\mathcal{L}_{\rm nd}+\lambda_{\rm sm}\mathcal{L}_{\rm sm}+\lambda_{\rm m}\mathcal{L}_{\rm m}+\lambda_{\rm\delta n}\mathcal{L}_{\rm\delta n},$
where $\mathcal{L}_{\rm c}$ is the color supervision between the rendered image and ground truth as in 3DGS, $\mathcal{L}_{\rm nd}$ is the supervision between the normal from Gaussian and from the estimated depth proposed in Gaussian Shader (Jiang et al., 2024), $\mathcal{L}_{\rm sm}$ is the smoothness loss for BRDF parameters, $\mathcal{L}_{\rm m}$ is the mask loss, and $\mathcal{L}_{\rm\delta n}$ is the delta normal regularization in Gaussian Shader. More details are shown in Sec. 5.
4.2. SDF-aided Gaussian splatting
Thanks to the deferred Gaussian splatting, the blurry issue for reflective object relighting has been addressed. However, the geometry established by Gaussians exhibits erroneous surfaces and floaters, as it tends to overfit the appearance under training light conditions. Hence, more constraints on the geometry are needed. Inspired by NeRO and TensoSDF, we introduce SDF into our framework, as it has shown to be a good prior for geometry regularization. The key is leveraging the capability of SDF while avoiding the expensive rendering. For this, we first choose a performance-friendly backbone for SDF – TensoSDF. Then, we propose geometry-only supervision between Gaussians and SDF, or so-called mutual supervision, together with the color supervision between Gaussians renderings and the input images, without any SDF rendering. On the one hand, as Gaussians build the link between rendered images and input images, they can supervise the training of SDF. On the other hand, the SDF can also constrain the Gaussians.
Mutual supervision
Even with a performance-friendly SDF backbone, TensoSDF rendering is still expensive, due to the complex color network required to faithfully represent the specular highlights. Therefore, we propose to remove the SDF rendering and connect the SDF and Gaussians with geometry attributes.
Specifically, we pose supervision on the depth and normal of the SDF (Eqn. (3)) from Gaussians (Eqn. (2)):
(8) $\begin{split}\mathcal{L}_{\rm gs2sdf}=|\textbf{sg}[D_{\rm gs}]-D_{\rm sdf}|+(1-<\textbf{sg}[\textbf{n}_{\rm gs}],\textbf{n}_{\rm sdf}>),\end{split}$
where $D_{\rm gs}$ and $D_{\rm sdf}$ are the depth from Gaussians and SDF, respectively, $\textbf{n}_{\rm gs}$ and $\textbf{n}_{\rm sdf}$ are normals from Gaussians and SDF respectively, and $\textbf{sg}[\cdot]$ is the stop-gradient operation.
Besides the above loss, we also utilize the Eikonal loss $\mathcal{L}_{\rm eik}$, Hessian loss $\mathcal{L}_{\rm hes}$, mask loss $\mathcal{L}_{\rm m}$, and total variance loss $\mathcal{L}_{\rm tv}$ to regularize the training of SDF, following common practice (Li et al., 2024; Gropp et al., 2020; Zhang et al., 2022c). The final loss for SDF is
(9) $\mathcal{L}_{\rm sdf}=\lambda_{\rm gs2sdf}\mathcal{L}_{\rm gs2sdf}+\lambda_{\rm eik}\mathcal{L}_{\rm eik}+\lambda_{\rm hes}\mathcal{L}_{\rm hes}+\mathcal{L}_{\rm m}+\lambda_{\rm tv}\mathcal{L}_{\rm tv},$
where the $\lambda_{[\cdot]}$ is the corresponding coefficient to adjust the strength of regularization. Among these losses, our proposed $\mathcal{L}_{\rm gs2sdf}$ enables the optimization of SDF via the geometry of Gaussians, and the $\mathcal{L}_{\rm eik},\mathcal{L}_{\rm hes},\mathcal{L}_{\rm m},\mathcal{L}_{\rm tv}$ regularize the SDF to obtain smooth and reasonable geometry.
Then, we design another loss to update Gaussians with SDF priors:
(10) $\mathcal{L}_{\rm sdf2gs}=|D_{\rm gs}-\textbf{sg}[D_{\rm sdf}]|+(1-<\textbf{n}_{\rm gs},\textbf{sg}[\textbf{n}_{\rm sdf}]>).$
Finally, we obtain the mutual supervision $\mathcal{L}_{\rm gs2sdf}$ and $\mathcal{L}_{\rm sdf2gs}$, which bridges the SDF and Gaussian. Note that the stop-gradient operator is used to differ the strength of the supervisions of both sides.
Discussion.
In general, SDF is optimized via color supervision between volume-rendered images and input images. However, our loss $\mathcal{L}_{\rm gs2sdf}$ can optimize the SDF via depth and normal without color supervision, as the depth and the normal are derived from SDF, and thus the gradient from $\mathcal{L}_{\rm gs2sdf}$ can optimize the SDF. In this way, we can drop the heavy color rendering and optimize the SDF with only geometry attributes (i.e., depth and normal). With other widely used regularizations for SDF (e.g., Eikonal loss), we can obtain a smooth and reasonable geometry from SDF. The loss $\mathcal{L}_{\rm sdf2gs}$ regularizes the geometry of Gaussians in turn, so we can obtain a better decomposition between geometry and material from the Gaussians.
4.3. SDF-aware pruning
Figure 5. In this example, the floaters can still be shown in the testing view, even if the mask loss was applied during training. The main reason is that although the floaters are shown outside the mask region in the testing view, they are within the mask region in the training views. Therefore, they can not be masked out by the mask loss.
Figure 6. The illustration of SDF-aware pruning. We define a narrowing threshold, which is adjusted automatically around the zero-level set. The Gaussians out of the threshold will be pruned. This pruning operation ensures all Gaussians are near the surface and avoids the floaters.
With the deferred Gaussians supervised by SDF priors, the geometry becomes smoother. However, we observe some artifacts caused by the overfitting on specular highlights of reflective objects in training views, which leads to Gaussian outliers distant from the surface. These overfitted Gaussians are in the mask region of all training views, so applying the mask loss cannot remove such outliers in training, and we still observe the floaters in the test view, as shown in  Fig. 5. Besides, all the above losses supervise the alpha-blended results and lack constraints on individual Gaussians to handle the outliers. Therefore, we propose a Gaussian pruning strategy to enforce the Gaussians close to the zero-level set of SDF. Specifically, for each Gaussian, we check the SDF value at its center and discard the Gaussian if the SDF value is larger than a threshold $s_{\varepsilon}$, as shown in the Fig. 6. Since it is laborious to adjust the threshold per scene, we link it with the probability density function $\phi_{s}(s)$ in Eqn. (4), and the threshold $s_{\varepsilon}$ can be defined as $\phi_{s}(s_{\varepsilon})=p_{t}(s_{\varepsilon}>0)$ in the closed form:
(11) $s_{\varepsilon}=\gamma\log{(2p_{t})}-\gamma\log{\left(\gamma^{-1}-2p_{t}-\sqrt{\gamma^{-2}-4p_{t}\gamma^{-1}}\right)},$
where $p_{t}$ is an empirical hyper-parameter and set to $0.01$. In this way, the threshold is determined automatically and narrowed with the SDF in the optimization, so we avoid posing a constant too strict when the surface is under-reconstructed and prune the Gaussians too far away from the surface in the end.
5. Implementation details
This section presents the network structures and the training details.
Network structures.
The TensoSDF has a resolution of $400\times 400$ with feature channels set as $36$. The MLP decoder is a two-layer MLP whose width is 128. The Gaussian starts with 100K randomly initialized points.
Table 1. Losses used in our paper.
Name Apply on Description
$L_{\rm c}$ GS Color supervision for 3DGS
$L_{\rm nd}$ GS Normal Consistency loss from Gaussian Shader.
$L_{\rm sm}$ GS BRDF smoothness loss from Relightable Gaussian.
$L_{\rm\delta n}$ GS Normal regularization for Gaussian Shader.
$L_{\rm sdf2gs}$ GS The proposed depth/normal supervision for GS.
$L_{\rm gs2sdf}$ SDF The proposed depth/normal supervision for SDF.
$L_{\rm eik}$ SDF Eikonal loss from Gropp et al. (2020).
$L_{\rm hes}$ SDF Hession Loss from Zhang et al. (2022c).
$L_{\rm tv}$ SDF Total variance loss from TensoRF.
$L_{\rm m}$ GS/SDF Mask loss for both 3DGS and SDF.
[…]
NeRF-based Gaussian-based
MII TensoIR TensoSDF NeRO GShader GS-IR RelightGS Ours
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
Angel 16.24/.8236 10.24/.2238 20.40/.8969 16.21/.7819 17.49/.8336 15.64/.6126 16.65/.8013 20.81/.8775
Bell 17.41/.8594 10.11/.1018 29.91/.9767 31.19/.9794 19.01/.8804 12.61/.2807 16.15/.8391 24.49/.9267
Cat 17.68/.8521 9.10/.1644 26.12/.9354 28.42/.9579 16.00/.8642 18.04/.7907 17.49/.8503 26.28/.9421
Horse 20.98/.8997 10.42/.1931 27.18/.9567 25.56/.9437 22.49/.9262 17.40/.7270 20.63/.8832 23.31/.9376
Luyu 17.89/.8050 8.27/.2375 19.91/.8825 26.22/.9092 15.62/.8254 19.00/.7727 17.47/.8168 22.61/.8995
Potion 17.13/.8094 6.21/.0846 27.71/.9422 30.14/.9561 12.33/.7575 16.37/.7051 14.99/.7799 25.67/.9175
Tbell 16.54/.8262 7.47/.1609 23.33/.9404 25.45/.9607 14.42/.8007 14.35/.5419 15.99/.7965 22.80/.9180
Teapot 16.71/.8546 9.96/.2093 25.16/.9482 29.87/.9755 18.21/.8560 16.63/.7646 17.36/.8389 21.17/.8932
Mean 17.57/.8413 8.97/.1719 24.97/.9349 26.63/.9331 16.95/.8430 16.26/.6494 17.09/.8258 23.39/.9140
Training Time 4h 5h 6h 12h 0.5h 0.5h 1h 1.5h
Ren. Time (FPS) 1/30 1/60 1/4 1/4 50 214 1.5 208
Optimization details
The training of our framework includes three steps. Firstly, the deferred Gaussian splatting without SDF supervision is trained separately for 1K iterations with Eqn. (7) to obtain a coarse geometry. Then, we freeze the Gaussians and warm up the TensoSDF via the depth and normal from Gaussians Eqn. (9) for 3K iterations. Finally, we train the entire framework with a joint loss for 24K iterations:
(12) $\mathcal{L}_{\rm joint}=\mathcal{L}_{\rm gs}+\lambda_{\rm sdf2gs}\mathcal{L}_{\rm sdf2gs}.$
We provide a table of loss functions (see  Tab. 1) for better understanding and more details of the losses are in the supplementary material.
We train our model on an NVIDIA RTX 4090 and optimize it with Adam optimizer (Kingma and Ba, 2017). The learning rate for the TensoSDF (Li et al., 2024) grid and the MLP decoder is $1e^{-2}$ and $1e^{-3}$, respectively. The learning rate of Gaussian components follows the setting in 3DGS (Kerbl et al., 2023), and all attributes not in vanilla 3DGS have a learning rate of $2e^{-3}$. The initial resolution of TensoSDF is set to $128$ and increases to the final resolution $400\times 400$, which uses only 60% capacity of the original TensoSDF, whose final resolution is $512\times 512$. After training, the TensoSDF will be discarded, and only Gaussians will be used for relighting. The pruning threshold is set to $0.01$ empirically. The optimization procedure takes about 1.5 hours in total.
6. Results
In this section, we describe the evaluation setup (Sec. 6.1), and evaluate our relighting and NVS results (Sec. 6.2). Then, we conduct ablation studies on several main components (Sec. 5) and discuss limitations (Sec. 6.4).
6.1. Evaluation setup
Dataset
We evaluate our method on three synthetic datasets and one real-scene dataset. We utilize the TensoIR (Jin et al., 2023) synthetic and Glossy Blender dataset (Liu et al., 2023) to evaluate the performance of diffuse and specular materials, respectively. Besides, we use the Shiny Blender dataset to show the generalization ability of our method for diverse materials. For real scenes, we select some specular objects from the NeILF++(Zhang et al., 2023).
Methods for comparison
We select 7 representative methods for comparison. We choose MII (Zhang et al., 2022b), TensoIR (Jin et al., 2023), TensoSDF (Li et al., 2024), and NeRO (Liu et al., 2023) for NeRF-based methods and Gaussian Shader (Jiang et al., 2024), GS-IR (Liang et al., 2024), and Relightable Gaussian (Gao et al., 2024) for Gaussian-based methods. We trained these models based on their public codes and configurations. Note that we add the mask loss to train all the compared methods, for stabilizing training and fair comparison.
Metrics
We use the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) (Wang et al., 2004) to measure the quality of results for the comparison and ablation study in the main paper. We present the learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) in the supplementary materials. Due to the ambiguity between albedo and light, we normalize the relighting images to match the average colors of ground-truth images before computing the metrics, following NeRO. When the ground truth albedo is available, we rescale the relighting images based on the ground truth albedo and predicted albedo, following TensoIR. We record the mean training time and the frame per second (FPS) [^1] NeRO and TensoSDF use the Cycles Render Engine in Blender to provide relighting results, so the FPS depends on samples per pixel, which is 1024 following the NeRO setting. in the evaluation, which tests on a single RTX 4090.
6.2. Comparison with previous works
Relighting of reflective objects
We evaluate the relighting performance of specular materials on the Glossy Blender dataset. The quantitative measurements are shown in Tab. 2, and the relighting visualizations are placed in Fig. 10. Our method outperforms other Gaussian-based methods and achieves competitive results of the NeRF-based methods. We also show the decomposed maps in Fig. 11 and Fig. 12. Due to the SDF priors and the deferred Gaussian pipeline, our method can produce a smooth surface without losing details and predict reasonable BRDF parameters.
Relighting of diffuse objects
Although our design focuses on reflective surfaces, it can also show benefits for diffuse objects. We validate our method on the TensoIR dataset. The quantitative measurements are shown in Tab. 3. The comparison with NeRF-based methods and relighting visualizations are in the supplementary material. Our methods can outperform the existing Gaussian-based methods and some NeRF-based methods on the TensoIR dataset, which reveals the robustness under diverse materials.
Relighting on real data
For real scenes, we select some objects with specular highlights from NeILF++ (Zhang et al., 2023), and the results are shown in Fig. 13. As there is no available ground truth, we only show the reference training views and some results under novel light conditions. Our method produces reasonable relighting results under novel lighting. Besides, our method achieves a balance between quality and training speed, enabling reflective surface modeling with only 25% training time of TensoSDF and 13% of NeRO. Our method supports real-time rendering with 200+ FPS, much faster than all NeRF-based and some Gaussian-based methods.
Table 3. Relighting quality in terms of PSNR↑ and SSIM↑ on the TensoIR synthetic dataset. Numbers in bold indicate the best performance, and underscored numbers indicate the second best.
GShader GS-IR RelightGS Ours
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
Armad. 22.86/.9280 27.65/.9078 30.76/.9526 31.33/.9593
Ficus 24.61/.9491 23.63/.8662 27.23/.9637 26.28/.9542
Hotdog 17.45/.8838 21.51/.8853 24.59/.9162 25.21/.9307
Lego 13.41/.7904 22.88/.8342 22.49/.8682 25.46/.9083
Mean 19.58/.8878 23.92/.8734 26.27/.9252 27.07/.9381
Table 4. NVS quality in terms of PSNR↑ and SSIM↑ on the Glossy Blender dataset. Numbers in bold indicate the best performance, and underscored numbers indicate the second best.
GS-IR RelightGS GShader Ours
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
Angel 20.77/.6935 23.12/.7841 27.14/.9226 29.32/.9445
Bell 19.05/.4895 24.38/.8952 30.00/.9409 31.53/.9694
Cat 27.91/.9075 29.73/.9487 31.25/.9604 31.72/.9672
Horse 20.85/.6835 22.88/.7769 26.03/.9314 27.09/.9479
Luyu 24.96/.8637 25.97/.8986 27.35/.9175 28.53/.9383
Potion 25.81/.8217 29.03/.9277 29.53/.9357 30.51/.9503
Tbell 20.72/.6193 23.03/.8878 23.86/.9031 29.48/.9648
Teapot 20.56/.8427 20.82/.8719 23.56/.8986 26.41/.9468
Mean 22.58/.6322 24.87/.8739 27.34/.9263 29.32/.9537
[…]
GS-IR RelightGS GShader Ours
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
Ball 18.30/.7584 21.39/.9047 30.40/.9623 35.50/.9849
Car 25.30/.8867 26.59/.9267 28.39/.9388 30.52/.9638
Coffee 30.72/.9463 32.57/.9710 30.79/.9690 29.64/.9568
Helmet 25.08/.9018 26.95/.9469 28.78/.9549 32.62/.9738
Teapot 38.21/.9900 43.86/.9963 43.35/.9957 43.88/.9964
Toaster 18.66/.7418 20.07/.8745 23.95/.9130 25.89/.9379
Mean 26.05/.8708 28.57/.9367 30.94/.9556 33.01/.9689
Figure 7. Ablation of several key components in our method, including the deferred shading, SDF supervision and pruning. The insets are shown in blue, and some floaters are shown in green.
NVS
We also present the NVS results on the Glossy Blender dataset and the Shiny Blender Dataset. The quantitative evaluation in terms of PSNR and SSIM is present in Tabs. 4 and 5. Our method can provide high-quality NVS results with sharper highlights, which are present in Figs. 14 and 15, revealing the wide applications of our method. Our method outperforms the existing Gaussian-based methods on most scenes. However, our method fails in the Coffee scene, as the deferred pipeline assumes an opaque surface, and the liquid in the Coffee does not meet the assumption. Besides, we present a comparison of normal estimation in Fig. 16 and the mean angular error (MAE) in Tab. 6. Thanks to the SDF priors, our method can provide robust normal estimation, while the results of other Gaussian-based methods are overfitted or noisy.
[…]
GS-IR RelightGS GShader Ours
Ball 25.79 22.44 7.03 0.92
Car 28.31 26.02 14.05 11.98
Coffee 15.38 13.39 14.93 12.24
Helmet 25.58 19.63 9.33 4.10
Teapot 15.35 9.21 7.17 5.88
Toaster 33.51 28.17 13.08 8.24
Mean 23.99 19.81 10.93 7.23
6.3. Ablation study
We conduct ablation studies on the key components of our model on the Glossy Blender dataset. The mean relighting PSNR and SSIM are presented in the Tab. 8. We can see a consistent improvement when we employ new components to the model. We start from a baseline that first computes colors for Gaussians and alpha blends them as in 3DGS. As shown in the 1st column in Fig. 7, the baseline cannot render a reflective surface and provide the result with obvious artifacts due to the blending issue discussed in Sec. 4.1.
Deferred Gaussian pipeline
After applying the deferred Gaussian pipeline, this variant provides a more specular result. However, due to the erroneous normal, the surface reflects light from the wrong directions, shown in the 2nd column of Fig. 7. We obverse more artifacts from the deferred results, for the pixel color from deferred shading corresponds to one shading point on the estimated surface while the one from forward shading is a combination of rendered colors of multiple Gaussians. The deferred formulation leaves less freedom for individual Gaussians to overfit. Consequently, more Gaussian outliers are generated to overfit sharp highlights, causing additional artifacts. Nevertheless, the deferred shading is necessary for specular materials (Sec. 4.1), and its issue has been solved in our full model.
TensoSDF incorporation
Then, we incorporate the TensoSDF and utilize its priors to regularize the geometry. The model without TensoSDF is vulnerable to the local minima and overfits under the training light condition, leading to erroneous geometry estimation. After employing the TensoSDF (3rd column in Fig. 7), the model mitigates the ambiguity between material and geometry and predicts a reasonable normal. We also validate the effectiveness of our mutual supervision by comparing it to supervision by a pretrained TensoSDF. As shown in Tab. 7, although using the pretrained TensoSDF for supervision could improve the quality slightly, it needs 5 hours for TensoSDF training. In contrast, our method achieves a balance between performance and efficiency. Besides, our method preserves the details while SDF-based methods fail, as shown in Fig. 8 (see the cat whisker), revealing the benefits of our Gaussian representation.
Figure 8. Our method preserves the details that are lost in the SDF-based method (i.e., NeRO), revealing the benefits of the Gaussian representation.
Table 7. Comparison of the relighting quality between mutual supervision (Ours) and using a pretrained TensoSDF to supervise Gaussians unidirectionally on the Glossy Blender dataset. Despite the slight quality improvement, using the pretrained model leads to an extensive training time cost.
Angel Cat Horse
Ours 20.81 26.28 23.31
Pretrained 21.45 (+0.64/+5h) 26.37 (+0.09/+5h) 24.59 (+1.28/+5h)
SDF-aware pruning
Eventually, we apply the SDF-aware pruning operation to our model. The models without pruning are likely to overfit under training views, and we can observe some unnecessary Gaussians under test views, causing severe floaters. Note that the overfitted Gaussians are invisible and in the mask region in training views. Therefore, we still observe the floater in the testing views, even though we apply the mask loss during training. After applying pruning, all Gaussians are located near the surface at the end of the training, and the model provides relighting results without floater, as shown in the 4th column in Fig. 7.
Table 8. Ablation study of three key components on the Glossy Blender dataset. “Def.” means the deferred Gaussian pipeline, “Inc.” means the incorporation of SDF and Gaussian, and “Pru.” means the pruning operation.
Components Scene
Def. Inc. Pru. Horse Cat Angel
PSNR/SSIM PSNR/SSIM PSNR/SSIM
20.33/0.9012 24.18/0.9253 18.86/0.8562
20.59/0.9089 24.87/0.9323 18.86/0.8574
23.12/0.9357 26.22/0.9417 20.48/0.8793
23.31/0.9376 26.28/0.9421 20.81/0.8775
6.4. Discussion and limitations
Our model improves the relighting quality compared to other Gaussian-based methods. However, some issues remain to be solved. The SDF representation used in our framework is used as a prior, and it can not be used for exporting a faithful mesh due to its low resolution. Besides, we obverse the split-sum approximation in DiffRast (Laine et al., 2020) causes blurry renderings for extremely low-roughness surfaces. We render a mirror-like plane using the Cycle path-tracing renderer in Blender and the split-sum approximation in the DiffRast. The comparison is in Fig. 9, and the split-sum result is blurry. We believe replacing it with a more accurate rendering method will improve the relighting quality. Last, we only consider direct lighting in our framework, and introducing indirect illumination will benefit the relighting quality. We leave it for future work.
Figure 9. The comparison between the result using Cycle path-tracing renderer in Blender and the one using split-sum approximation in DiffRast. The split-sum result is blurry with extremely low roughness.
Figure 10. Relighting results on the Glossy Blender Dataset. Relightable Gaussian, GS-IR, TensoIR, and MII fail to model specular highlights. Gaussian Shader provides specular relighting results with obvious artifacts. Our method, along with NeRO and TensoSDF, provides high-quality results, while our method only uses 25% training time of TensoSDF and enables real-time relighting. We present the metrics per scene (PSNR/SSIM).
Figure 11. Decomposed maps of our method on the Glossy Blender Dataset. Our method can provide a reasonable decomposition for reflective surfaces. The normal is smooth but maintains details (see the horse mane in the 2nd row). There is no material ground truth, so we only provide the qualitative visualization.
Figure 12. Normal comparison among Gaussian-based methods on the Glossy Blender Dataset. Our method can reconstruct high-quality normal for relighting, while others overfit under the training scene and provide erroneous normal.
Figure 13. Relighting results on real data from NeILF++. Relightable Gaussian and GS-IR fail to reconstruct the reflective objects in real scenes. Gaussian Shader cannot relight the specular highlights faithfully, while our method can provide reasonable relighting results.
Figure 14. NVS results on the Glossy dataset. Our method provides photo-realistic NVS results for reflective objects, while other Gaussian-based methods fail.
Figure 15. NVS results of Gaussian-based methods on the Shiny Blender dataset. Our method preserves correct highlights, while others fail to render sharp highlights.
Figure 16. Normal comparison with Gaussian-based methods on the Shiny Blender dataset. Our method provides robust normal estimation, while other results are noisy or overfitted.
7. Conclusion
In this paper, we present a novel framework for real-time reflective object inverse rendering. We design an SDF-aided Gaussian splatting framework, using the mutual supervision of the depth and normal between deferred Gaussians and SDF to improve the quality of geometry from Gaussians. Besides, we propose an SDF-aware pruning strategy with an automatically adjusted threshold, regularizing the position of Gaussian and avoiding the floater artifact. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods without losing efficiency. It is competitive with the NeRF-based methods in terms of quality, with much less training and rendering time. In future work, extending our framework to complex scenarios with multiples objects is a potential direction.